# PointDreamer: Zero-shot 3D Textured Mesh Reconstruction from Colored Point Cloud

### v2

@@ Abstract
Reconstructing textured meshes from colored point clouds is an important but challenging task. Most existing methods yield blurry-looking textures or rely on 3D training data that are hard to acquire. Regarding this, we propose PointDreamer, a novel framework for textured mesh reconstruction from colored point cloud via diffusion-based 2D inpainting. Specifically, we first reconstruct an untextured mesh. Next, we project the input point cloud into 2D space to generate sparse multi-view images, and then inpaint empty pixels utilizing a pre-trained 2D diffusion model. After that, we unproject the colors of the inpainted dense images onto the untextured mesh, thus obtaining the final textured mesh. This project-inpaint-unproject pipeline bridges the gap between 3D point clouds and 2D diffusion models for the first time. Thanks to the powerful 2D diffusion model pre-trained on extensive 2D data, PointDreamer reconstructs clear, high-quality textures with high robustness to sparse or noisy input. Also, it’s zero-shot requiring no extra training. In addition, we design Non-Border-First unprojection strategy to address the border-area inconsistency issue, which is less explored but commonly-occurred in methods that generate 3D textures from multiview images. Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets show the SoTA performance of PointDreamer, by significantly outperforming baseline methods with 30% improvement in LPIPS score (from 0.118 to 0.068). Code at: https://github.com/YuQiao0303/PointDreamer.
@@ 1 Introduction
3D point clouds and meshes are two of the most fundamental 3D modalities in computer vision, graphics, and robotics. Reconstructing 3D textured meshes from colored point clouds (abbr. colored-PC-to-mesh) is a pivotal task, with extensive applications like digital twin [1], metaverse [2], cultural heritage preservation [3], etc. It requires the accurate reconstruction of both geometry and colors, to transform sparse and unstructured data into coherent and visually appealing models.
While recent methods [4, 5, 6] have achieved relatively satisfactory geometry reconstruction quality, color reconstruction remains challenging, especially for low-density point clouds. Unlike images that provide dense pixel-level representations, the inherent sparse nature of point clouds makes it difficult to reconstruct dense colors. Regarding this, existing methods either blend input points’ colors by interpolation [7] or overfitting [8, 9], or train color prediction networks [10, 11] with 3D datasets. However, they often yield blurring-looking textures; see SPR [7], NKSR [9], and Texture Field [12, 10] in Figure 1. Also, 3D training data are notoriously challenging to acquire.
@@
However, it is not trivial to implement this idea: (1) Existing 2D-diffusion-for-3D methods typically focus on 3D generation conditioned on text [15, 16] or images [17, 18], leaving point-cloud-based reconstruction unexplored. (2) Existing 2D-diffusion-for-3D methods often yield blurry appearances [19] by predicting colors in 3D space as a texture [20, 16]/radiance [21, 22, 23] field or 3D Gaussians [19, 18]. How to leverage 2D reference images to generate high-quality textures requires more attention. (3) Intermediate 2D multiview images generated by diffusion models often have inconsistencies with each other or with the reconstructed 3D geometry at occlusion border areas. When one part of an object occludes another, the generated 2D images and 3D geometry should predict highly consistent occlusion borders. Failing to do so would cause artifacts; see “Ours w/o/ NBF” in Figure 1, with significant artifacts near the border of the pillow and the chair.
@@
Putting everything together, we propose PointDreamer, a novel zero-shot framework to reconstruct high-quality textured meshes from colored point clouds, as shown in Figure 1. First, we extract geometry (untextured mesh) from the point cloud by an existing method POCO [4]. Second, we project the point cloud into 2D space from a fixed set of viewpoints, producing multiview sparse images. Third, we inpaint the empty pixels with an off-the-shelf 2D diffusion model, forming dense images. Finally, we directly unproject the 2D images onto the mesh. Unlike most existing methods that predict colors in 3D [10, 11] or UV spaces [8], directly adopting colors of 2D diffusion models’ output high-quality 2D images leads to clear textures. In particular, our designed NBF unprojection strategy effectively avoids artifacts caused by inconsistencies in occlusion border regions.
@@
  • We propose PointDreamer, a zero-shot SoTA framework that fills the gap in the less-explored but important task of 3D textured mesh reconstruction from colored point cloud, pioneering in utilizing 2D diffusion models for 3D point cloud reconstruction.
  • PointDreamer reconstructs high-quality and clear-looking textures for meshes, avoiding the common blurring issue. PointDreamer is also zero-shot. While requiring no extra training, it shows high generalization capability: robust to noisy, sparse, or even incomplete input points, benefiting from the power of 2D diffusion models.
  • We design Non-Border-First (NBF) Unprojection, a novel unprojection strategy to apply colors of 2D multiview images onto 3D meshes. It effectively avoids artifacts caused by inconsistent prediction near occlusion borders in any multiview-images-to-3D methods.
@@ 2.1 Textured Mesh Reconstruction from Colored Point Cloud
2.2 2D Diffusion Models for 3D Generation
To capitalize on the abundant availability of 2D data and the high performance of 2D diffusion models, researchers seek to “lift” results of 2D diffusion models to 3D. Most of these works optimize or learn a NeRF [22, 23], a DMTet with texture field [29, 16], or 3D Gaussianss by Score Distillation Sampling [13] or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model. View-conditioned diffusion models [22] or multi-view diffusion models [30, 14, 31] are developed by fine-tuning existing diffusion models for generating multi-view-consistent images. Most of these works focus on 3D generation from text or images instead of point clouds. They tend to generate blurry textures, and may suffer from artifacts in occlusion border areas.
@@ 3 Method
Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training. It takes only about 61s and 30s for PointDreamer’s DDNM and nearest/linear interpolation variants (see Figure 7 (b)) to reconstruct a shape of 30k points on an NVIDIA A100 GPU, compared to other 2D-diffusion-for-3D methods such as Zero-1-to-3 [22] (~20 m), DreamGaussin [19] (~2 m), Wonder3D [32] (2~3 m), LGM [18] (~65 s), etc.
Figure 2: Illustration of hidden point removal, without which the results would be messed up.
Figure 3: Top: Illustration of input and output of the unprojection module, which unprojects colors of 2D multi-view images to a 3D mesh. Middle Left: An naive unprojection implementation. For each 3D point $\mathbf{p}$, from all views where $\mathbf{p}$ is visible, the highest-direction-priority view is chosen to paint $\mathbf{p}$. Example point $\mathbf{p}$’s corresponding pixel in the selected view $a$ is with incorrect color since it’s near an occlusion border, producing an artifact. Bottom Right: our proposed Non-Border-First Unprojection. Observing the correspondence between the occlusion border in 2D image space and the visibility border in UV space (both marked by sky-blue dashed lines), we prioritize non-border areas in UV space to avoid such artifacts.
@@ 3.2 Projection: 3D to 2D
Directly projecting the whole point cloud into 2D space would incorporate points from occluded parts that should not be visible from the given view. This would mess up the inpainting results and subsequent mesh textures; see Figure 2 for an example. Therefore, before generating the sparse images, we process the input point cloud by the “Hidden Point Removal” operator [41], which works by transforming the input and extracting the points that reside on its convex hull. Besides, we compare the depth values of points and the extracted untextured mesh to further remove some invisible points. With pre-set camera parameters of $K$ viewpoints ($K=8$ in our implementation) and associated visible point clouds, we conduct camera transformation to these points to get their corresponding pixel coordinates in 2D image space. These pixels are painted according to their associated 3D points’ colors, producing $K$ sparse images. We compare different $K$ values in our supplementary file.
Algorithm 1 PyTorch-like pseudocode for Non-Border-First (NBF) strategy
1# K: number of views
2# T: texture atlas
3# I: multiview images
4# I_dep: [K, I_res, I_res], multi-view depth maps obtained from geometry
5# T_dep: [K, T_res, T_res], T’s pixels’ corresponding depths in each view
6# Tuv2Iab: [T_res, T_res, 2]: Mapping T’s pixel coordinate to I’s
7# chart_mask: [T_res, T_res]: Does T(u,v) belong to a chart (in foreground)?
8# T_normal: [T_res, T_res, 3]: T’s pixels’ corresponding 3D point’ normal directions
9# view_dir: [K, 3]: camera view directions
10# I_rgb: [K, I_res, I_res, 3], multi-view RGB images
11T = torch.zeros(T_res, T_res, 3)
12vis_T = calc_vis(I_dep, T_dep, Tuv2Iab) # [K, T_res, T_res], bool
13vis_T_edges = detect_edge(vis_T) # [K, T_res, T_res], bool
14chart_edges = detect_edge(chart_mask) # [T_res, T_res], bool
15border_edges = vis_T_edges - chart_edges.unsqueeze(0).repeat(K)
16border_areas = dilate(border_edges, kernel) # [K, T_res, T_res], bool
17cand_view_T_map = non_border_areas = ~border_areas # [K, T_res, T_res]
18Tuv_cand_view_num = non_border_areas.sum(0) # [T_res, T_res], int
19cand_view_T_map[:, Tuv_cand_view_num==0] = vis_T[:, Tuv_cand_view_num==0]
20dir_priority = calc_priority(T_normal, view_dir) # [K, T_res, T_res]
21dir_priority[~cand_view_T_map] -= 1000 # Non-Border-First
22uv_view_id = dir_priority.argmax(0) # [T_res, T_res]
23for u in range(0, T_res):
24 for v in range(0, T_res):
25 T[u, v] = I_rgb[uv_view_id[u, v], Tuv2Iab[u, v, 0], Tuv2Iab[u, v, 1]]
@@ 3.4 Unprojection: 2D to 3D
The top of Figure 3 illustrates the basic idea of the unprojection module. Specifically, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$. We follow existing works [42, 32] to conduct UV mapping by Xatlas [43]. A texture atlas is an RGB image containing different charts. Each chart is a continuous area in UV space that corresponds to a continuous segment of $\mathcal{M}$ in 3D space. Each pixel $\mathcal{T}[u,v]$ of a chart corresponds to a surface point $\mathbf{p}$ of $\mathcal{M}$. Therefore, our goal is to assign a color for each chart pixel $\mathcal{T}[u,v]$. Given a set of multi-view posed images $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$, $\mathbf{p}$ can be seen in zero, one, or more images. If $\mathbf{p}$ is visible in $I_{k}$, we denote $I_{k}$’s corresponding pixel as $I_{k}[i_{k},j_{k}]$. See the top of Figure 3 as an example, where a 3D point $\mathbf{p}$ is visible in two views $a$ and $b$, and we mark it with a red circle together with its corresponding multi-view image pixels $I_{a}[i_{a},j_{a}]$, $I_{b}[i_{b},j_{b}]$, and texture atlas pixel $\mathcal{T}[u,v]$.
@@ 3.4.1 Naive Unprojection: Only Use Direction Priority
An intuitive idea to pick the best view for $\mathbf{p}$ or $\mathcal{T}[u,v]$ is by direction priority, as shown in the middle left part of Figure 3. Specifically, for each view $k$ where point $\mathbf{p}$ can be seen, we calculate the direction priority score $s_{\mathbf{p}}(k)$, defined as the cosine similarity of $\mathbf{p}$’s normal direction $\mathbf{n}$ and view $k$’s camera direction $\mathbf{d}_{k}$. The view with the highest direction priority is chosen to paint $\mathbf{p}$ and $\mathcal{T}[u,v]$. For example, in the middle left part of Figure 3, $\mathbf{p}$ is painted to gray according to $I_{a}[i_{a},j_{a}]$ since view $a$ has a higher direction priority. Also, if there are a few points invisible from any view, we can still assign a view by direction priority.
While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the middle left part of Figure 3, where a gray artifact is visible in the reconstructed textured mesh that should have been brown. Through analysis, such artifacts appear in border areas: the area near the border of occluded and occluding part of the shape when viewing from view $k$. For example, in view $a$, the brown seat’s bottom side is occluded by the gray bottom of the chair. The area near the edge of the brown and gray parts is a border area. In such areas, the inpainting module and the geometry extraction module may inconsistently delineate the boundary separating the two parts. For example, the gray bottom is larger in the inpainted image $I_{a}$ than of the reconstructed mesh, thus resulting in an artifact.
@@ 3.4.2 Optimization-based Unprojection
Instead of selecting one single view by direction priority, 3D generation method DreamGaussian [19] adopts an optimization-based strategy to get the textured mesh from multi-view images. The idea is to optimize the texture atlas by a per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh. In this way, the optimization process kinds of fuses corresponding colors in different views instead of selecting the best one. However, it cannot ignore the inconsistent colors in the border areas, so the artifacts can only be reduced instead of eliminated. This inspires us to design a non-border-first strategy to address this issue.

### v3

@@ Abstract
Faithfully reconstructing textured meshes is crucial for many applications. Compared to text or image modalities, leveraging 3D colored point clouds as input (colored-PC-to-mesh) offers inherent advantages in comprehensively and precisely replicating the target object’s 360° characteristics. While most existing colored-PC-to-mesh methods suffer from blurry textures or require hard-to-acquire 3D training data, we propose PointDreamer, a novel framework that harnesses 2D diffusion prior for superior texture quality. Crucially, unlike prior 2D-diffusion-for-3D works driven by text or image inputs, PointDreamer successfully adapts 2D diffusion models to 3D point cloud data by a novel project-inpaint-unproject pipeline. Specifically, it first projects the point cloud into sparse 2D images and then performs diffusion-based inpainting. After that, diverging from most existing 3D reconstruction or generation approaches that predict texture in 3D/UV space thus often yielding blurry texture, PointDreamer achieves high-quality texture by directly unprojecting the inpainted 2D images to the 3D mesh. Furthermore, we identify for the first time a typical kind of unprojection artifact appearing in occlusion borders, which is common in other multiview-image-to-3D pipelines but less-explored. To address this, we propose a novel solution named the Non-Border-First (NBF) unprojection strategy. Extensive qualitative and quantitative experiments on various synthetic and real-scanned datasets demonstrate that PointDreamer, though zero-shot, exhibits SoTA performance ( 30% improvement on LPIPS score from 0.118 to 0.068), and is robust to noisy, sparse, or even incomplete input data. Code at: https://github.com/YuQiao0303/PointDreamer.
@@ 1 Introduction
Faithful 3D mesh reconstruction is a fundamental task in computer vision and graphics. It aims to produce accurate digital recreation of the full geometry and appearance of real-world objects. While 3D generation [1, 2, 3] or mesh texturing models [4, 5, 6] from text or image inputs have seen remarkable advances, they produce results that exhibit only coarse semantic alignment or viewpoint-specific consistency compared to the target object. As a result, they cannot achieve 360° physically accurate reconstructions with high fidelity. In contrast, 3D colored point clouds provide direct measurements of an object’s surface geometry and colors from all perspectives. This essential feature theoretically provides a unique foundation for faithful reconstruction. Consequently, reconstructing coherent and visually appealing 3D textured meshes from sparse and unstructured colored point clouds (abbr. colored-PC-to-mesh) is a pivotal task, with extensive applications like digital twin [7], metaverse [8], cultural heritage preservation [9], etc.
Different from 3D generation based on text or images, point cloud reconstruction faces its own challenges. While recent point cloud reconstruction methods [10, 11, 12] have achieved relatively satisfactory geometry reconstruction quality, color reconstruction remains challenging, especially for low-density point clouds. Unlike images that provide dense pixel-level representations, the inherent sparse nature of point clouds makes it difficult to reconstruct dense colors. Regarding this, existing methods either blend input points’ colors by interpolation [13] or overfitting [14, 15], or train color prediction networks [16, 17] with 3D datasets. However, they often yield blurring-looking textures; see SPR [13], NKSR [15], and Texture Field [18, 16] in Figure 1. Also, 3D training data are notoriously challenging to acquire.
@@
However, it is not trivial to implement this idea: (1) Existing 2D diffusion models [20, 21] are predominantly developed to be conditioned on text or image inputs, fundamentally incompatible with point cloud data. Consequently, existing 2D-diffusion-based 3D generation methods are driven typically by text [22, 23] or images [3, 24], leaving point-cloud-based reconstruction under-explored. (2) Existing 2D-diffusion-for-3D methods often yield blurry appearances [25] by predicting colors in 3D space as a texture field [26, 23], a radiance field [27, 28, 29] or 3D Gaussians [25, 24]. How to leverage 2D reference images to generate high-quality textures requires more attention. (3) Intermediate 2D multiview images generated by diffusion models often have inconsistencies with each other or with the reconstructed 3D geometry at occlusion border areas. When one part of an object occludes another, the generated 2D images and 3D geometry should predict highly consistent occlusion borders. Failing to do so would cause artifacts; see “Ours w/o/ NBF” in Figure 1, with significant artifacts near the border of the pillow and the chair.
@@
Putting everything together, we propose PointDreamer, a novel zero-shot framework to reconstruct high-quality textured meshes from colored point clouds, as shown in Figure 1. First, we extract geometry (untextured mesh) from the point cloud by an existing method POCO [10]. Second, we project the point cloud into 2D space from a fixed set of viewpoints, producing multiview sparse images. Third, we inpaint the empty pixels with an off-the-shelf 2D diffusion model, forming dense images. Finally, we directly unproject the 2D images onto the mesh. Unlike most existing methods that predict colors in 3D [16, 17] or UV spaces [14], directly adopting the colors of 2D diffusion models’ output high-quality 2D images leads to clear textures. In particular, our designed NBF unprojection strategy effectively avoids artifacts caused by inconsistencies in occlusion border regions.
@@
  • We propose PointDreamer, a SoTA framework for 3D textured mesh reconstruction from colored point cloud, with multiple advantages: high-quality texture, zero-shot, and robust to noisy, sparse, or even incomplete input data.
  • Existing 2D diffusion models are predominantly developed for text or image inputs, thus confining existing 2D-diffusion-for-3D methods to these inputs too. Unlike them, PointDreamer resolves the intrinsic incompatibility between 2D diffusion and colored 3D point clouds, by a novel project-inpaint-unproject pipeline.
  • Many existing 3D reconstruction or generation methods yield blurry texture by predicting colors in 3D or UV space. In contrast, PointDreamer achieves high-quality texture by directly unprojecting predicted 2D images to 3D.
  • We identify for the first time a common but less-explored artifact in multiview-image-to-3D pipelines caused by inconsistent prediction near occlusion borders. We propose a novel strategy named Non-Border-First (NBF) unprojection to effectively address this issue.
@@ 2.1 Textured Mesh Reconstruction from Colored Point Cloud
2.2 2D Diffusion for 3D Generation and Texturing
With the success of 2D diffusion models, researchers have explored utilizing them for 3D generation and 3D mesh texturing. Most of these generation works optimize or learn a NeRF [28, 29], a DMTet with texture field [2, 35, 23], or 3D Gaussianss by Score Distillation Sampling [1] or reconstruction loss to make the generated 3D scene similar to the result of a given 2D diffusion model. View-conditioned diffusion models [28] or multi-view diffusion models [36, 19, 37] are developed to produce multiview-consistent images for subsequent 3D generation or texturing. Later, Large Reconstruction Models [38, 39, 3] are proposed, demonstrating the effectiveness of directly using feed-forward networks for 3D generation. In addition, inpainting-based methods are popular in the field of generating texture for 3D meshes [4, 5, 6]. They use geometry-aware diffusion models to progressively inpaint the complete texture. However, conditioned on only text or a single image instead of a point cloud, the above methods align more closely with “generation” instead of “reconstruction”. In other words, they theoretically cannot faithfully reconstruct the target object; see Figure 2 for an example. Also, they may generate blurry textures as shown in “CRM” in Figure 2, and may suffer from artifacts in occlusion border areas.
Figure 2: Comparison with different 2D-diffusion-for-3D methods.
@@ 3 Method
Note that with the 2D diffusion priors, our method is zero-shot requiring no extra training. It takes only about 61s for PointDreamer to reconstruct a shape of 30k points on an NVIDIA A100 GPU, compared to other 2D-diffusion-for-3D methods such as Zero-1-to-3 [28] (~20 m), DreamGaussin [25] (~2 m), Wonder3D [40] (2~3 m), LGM [24] (~65 s), Texture [4] (~5 m), Text2Tex [5] (~15 m), Easi-Tex [6] (~6 m), and ProlificDreamer [2] (several hours).
Figure 3: Hidden point removal aims to avoid significant stippled artifacts
Figure 4: The unprojection module takes an untextured mesh and corresponding multiview images as input, to output the associated textured mesh.
Figure 5: The naive unprojection strategy chooses one best view for each texture atlas pixel by considering only the direction priority (similarity between the 3D point normal direction and the camera direction). This can lead to border-area artifacts. The optimization-based unprojection strategy uses multiview images as a reference to optimize the texture atlas with per-pixel MSE loss. However, the artifacts still exist.
Figure 6: Illustration of the proposed NBF unprojection strategy. By painting each view’s visible texture atlas, we can detect its visibility border areas, which correspond to occlusion borders in the corresponding view’s 2D image. Then paint the texture by prioritizing non-border areas.
@@ 3.2 Projection: 3D to 2D
Directly projecting the whole point cloud into 2D space would incorporate points from occluded parts that should not be visible from the given view. This would cause significant stippled artifacts in the inpainting results and subsequent mesh textures; see Figure 3 for an example. Therefore, before generating the sparse images, we process the input point cloud by the “Hidden Point Removal” operator [49], which works by transforming the input and extracting the points that reside on its convex hull. Besides, we compare the depth values of points and the extracted untextured mesh to further remove some invisible points. With pre-set camera parameters of $K$ viewpoints ($K=8$ in our implementation) and associated visible point clouds, we conduct camera transformation to these points to get their corresponding pixel coordinates in 2D image space. These pixels are painted according to their associated 3D points’ colors, producing $K$ sparse images. We compare different $K$ values in our supplementary file.
@@ 3.4 Unprojection: 2D to 3D
Figure 4 illustrates the basic idea of the unprojection module. Specifically, we represent mesh texture as a texture atlas $\mathcal{T}$ by applying UV mapping to a mesh $\mathcal{M}$. We follow existing works [50, 40] to conduct UV mapping by Xatlas [51]. A texture atlas is an RGB image containing different charts (e.g. ①, ② … in Figure 4). Each chart is a continuous area in UV space that corresponds to a continuous segment of $\mathcal{M}$ in 3D space. For example, in Figure 4, chart ① corresponds to the bottom side of the brown seat, and chart ② corresponds to the bottom side of the gray cone. Each pixel $\mathcal{T}[u,v]$ of a chart corresponds to a surface point $\mathbf{p}$ of $\mathcal{M}$. Therefore, our goal is to assign a color for each chart pixel $\mathcal{T}[u,v]$. Given a set of multi-view posed images $\mathbf{I}=\{I_{k}\}_{k=1}^{K}$, $\mathbf{p}$ can be seen in zero, one, or more images. If $\mathbf{p}$ is visible in $I_{k}$, we denote $I_{k}$’s corresponding pixel as $I_{k}[i_{k},j_{k}]$. See the top of Figure 4 as an example, where a 3D point $\mathbf{p}$ is visible in two views $a$ and $b$, and we mark it with a red circle together with its corresponding multi-view image pixels $I_{a}[i_{a},j_{a}]$, $I_{b}[i_{b},j_{b}]$, and texture atlas pixel $\mathcal{T}[u,v]$.
@@ 3.4.1 Naive Unprojection: Only Use Direction Priority
An intuitive idea to pick the best view for $\mathbf{p}$ or $\mathcal{T}[u,v]$ is by direction priority, as shown in ‘Naive Unprojection Strategy’ on top of Figure 5. Specifically, for each view $k$ where point $\mathbf{p}$ can be seen, we calculate the direction priority score $s_{\mathbf{p}}(k)$, defined as the cosine similarity of $\mathbf{p}$’s normal direction $\mathbf{n}$ and view $k$’s camera direction $\mathbf{d}_{k}$. The view with the highest direction priority is chosen to paint $\mathbf{p}$ and $\mathcal{T}[u,v]$. For example, in the top row of Figure 5, $\mathbf{p}$ is painted as gray according to $I_{a}[i_{a},j_{a}]$ since view $a$ has a higher direction priority. Also, if there are a few points invisible from any view, we can still assign a view by direction priority.
While this naive unprojection strategy serves adequately for most points, it produces artifacts under certain circumstances; see again the top row of Figure 5, where a gray artifact is visible in the reconstructed textured mesh that should have been brown. As shown in ‘Analysis’ in Figure 5, such artifacts appear in border areas: the area near the border of the occluded and occluding part of the shape when viewing from view $k$. For example, in view $a$, the brown seat’s bottom side is occluded by the gray bottom of the chair. The area near the edge of the brown and gray parts is a border area. In such areas, the inpainting module and the geometry extraction module may inconsistently delineate the boundary separating the two parts. For example, the gray bottom is larger in the inpainted image $I_{a}$ than of the reconstructed mesh, thus resulting in the artifact.
@@ 3.4.2 Optimization-based Unprojection
Instead of selecting one single view by direction priority, 3D generation method DreamGaussian [25] adopts an optimization-based strategy to get the textured mesh from multi-view images. As shown in ‘Optimization Unprojection Strategy’ of Figure 5, the idea is to optimize the texture atlas by a per-pixel loss between the inpainted multi-view images $\{I_{k}\}_{k=1}^{K}$ and the rendered images of the reconstructed textured mesh. In this way, the optimization process fuses corresponding colors in different views instead of selecting the best one. However, it cannot ignore the inconsistent colors in the border areas, so the artifacts can only be reduced instead of eliminated. This inspires us to design a non-border-first strategy to address this issue.